OpenStudy (anonymous):

find the quadratic equation that has the 2 given roots. (below)

OpenStudy (anonymous):

ill try and help

OpenStudy (anonymous):

emphasis on the try

OpenStudy (anonymous):

\[\frac{ 2+\sqrt{5} }{ 3 }\]\[\frac{ 2-\sqrt{5} }{ 3 }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so instead of doing the quadratic equation they just give you the roots of it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

that's pretty weird, im sorry i have no idea how to do that. copy your question then close the question then paste it again, maybe someone else who is smarter than me can help:p

OpenStudy (anonymous):

alright, well thanks for trying

OpenStudy (anonymous):

if your in algebra one and you need help with a difffrent question i can help, you can do this : @socksarelame and it will tell me you need help

OpenStudy (anonymous):

Algebra 2 actually, but I appreciate your help :)